# QCS:Feature Refining from Quadruplet Cross Similarity for Facial Expression Recognition

### v1

@@ Abstract
On facial expression datasets with complex and numerous feature types, where the significance and dominance of labeled features are difficult to predict, facial expression recognition(FER) encounters the challenges of inter-class similarity and intra-class variances, making it difficult to mine effective features. We aim to solely leverage the feature similarity among facial samples to address this. We introduce the Cross Similarity Attention (CSA), an input-output position-sensitive attention mechanism that harnesses feature similarity across different images to compute the corresponding global spatial attention. Based on this, we propose a four-branch circular framework, called Quadruplet Cross Similarity (QCS), to extract discriminative features from the same class and eliminate redundant ones from different classes synchronously to refine cleaner features. The symmetry of the network ensures balanced and stable training and reduces the amount of CSA interaction matrix. Contrastive residual distillation is utilized to transfer the information learned in the cross module back to the base network. The cross-attention module exists during training, and only one base branch is retained during inference. our proposed QCS model outperforms state-of-the-art methods on several popular FER datasets, without requiring additional landmark information or other extra training data. The code is available at https://github.com/birdwcp/QCS.
@@ Introduction
However, FER is still a complex task facing many difficulties such as Inter-class similarity and intra-class variances. Faces with varying expressions retain the common features that constitute a human face, yet same expressions can vary widely in pose, identity, illumination, etc. Traditional object classification methods struggle with complex face datasets due to assumptions of labeled feature dominance. Dividing datasets by labels for training can result in the learning of unlabeled redundant features if they’re non-randomly distributed, impacting classification. Fig. 1 illustrates an extreme case of this impact. In reality, interference features are varied, numerous, and unpredictable, exerting varying degrees of influence.
Figure 1: Assuming that unlabeled gender features and labeled expression features both exhibit similar distributions in the training set, leading to the non-dominance of significant expression features. It is difficult to explain whether the trained model is an expression classifier or of another type.
@@
How to extract discriminative features and how to eliminate the interference of redundant features are the core issues of FER. Using self-attention, some deep learning methods (Wang et al. 2020; Xue, Wang, and Guo 2021; Xue et al. 2022; Ma, Sun, and Li 2021) focus on key features, enhancing model discrimination. However, model performance hinges on the amount and quality of training data. Some methods (Xie, Hu, and Wu 2019; Zhang et al. 2021; Cai et al. 2021; Ruan et al. 2021, 2022) study separating redundant from expression features using GANs or transformers. They extract expression information via decomposition and reconstruction, but often need redundant feature labels, such as age, gender, race, or rely on identity recognition modules. Additionally, interference decoupling FER (Li, Deng, and Du 2017) relies on one or several interference feature labels and can only address limited and specific interference factors. However, interference features are often unclear in type and quantity, and these methods that rely on labeled data often fail to achieve good performance when the types, quantities, and quality of labeled data are limited. Furthermore, when the expression feature labels are ambiguous (She et al. 2021; Zhang, Wang, and Deng 2021), it is unrealistic to clearly label all interference features.
Unlike the aforementioned supervised methods, we introduce a straightforward yet potent approach. It harnesses similarities among same-class image pairs to extract discriminative label features and employs similarities in cross-class image pairs to separate unlabeled, redundant interference features, to alleviate the problem of relying on annotated features to be dominant in the training set.
Firstly, We propose Cross Similarity Attention (CSA) tailored for mining fine-grained feature similarity between different images. It distinguishes from NLP Transformer’s cross-attention(Vaswani et al. 2017) for language correlations and Vision Transformer’s self-attention(Dosovitskiy et al. 2021) for intra-image correlations. CSA calculates Euclidean distance similarity between fine-grained features of images, focusing solely on the importance of the features in the spatial dimension of the queried image(key). This compensates for ViT’s matrix multiplication-based attention, which lacks input-output positions correlation in cross-image similarity.
@@
By leveraging CSA for interaction module construction, we enhance the distinct shared features between images within the same class. Inspired by triplet loss (Schroff, Kalenichenko, and Philbin 2015), we apply the interaction module to extract and diminish similar features across different class image pairs, thereby separating redundant and interfering features. To ensure equal feature refinement across input branches, we introduce a four-branch configuration, with each class contributing two images. This forms a centrally symmetric closed-loop network utilizing four sets of interaction modules for stable joint training. This approach simultaneously minimizes intra-class feature differences and maximizes inter-class differences, mimicking the effect of triplet loss at the fine-grained feature level. During inference, only one image is input, requiring the removal of the multi-branch interaction module, retaining a single base network. To effectively integrate the refined information from cross modules, we introduce residual connections that minimize gradient conflicts in feature maps, creating a self-distillation-like structure.
@@
  • We introduce Cross Similarity Attention (CSA), a tailored mechanism for mining fine-grained feature similarities across different images, utilizing a spatially dimensioned and input-output position-sensitive attention approach.
  • We propose Quadruplet Cross Similarity (QCS), a four-branch, centrally symmetric closed-loop joint training framework based on CSA. It refines features by mining discriminative features within the same class and separating interfering features across different classes synchronously.
  • Our method achieves state-of-the-art performance on several FER datasets without requiring additional landmark information or pre-training on FER external datasets.
Figure 2: The framework of Quadruplet Cross Similarity Network. Joint training is performed on 4 classifiers ${{Cls}_{base}}$ based on a weight-shared backbone and 4 classifiers ${{Cls}_{cross}}$ enhanced by cross-module, with only the red branch retained during inference. Anchor and pos are in the same category, so are neg and neg2. The matrix ${SD}$ performs attention refinement on matrix ${S}$ and matrix ${D}$ by rows or columns.
@@
Transfer(Xue, Wang, and Guo 2021) is the first attempt to apply Vision Transformer (ViT) to adaptively characterize the relations between different facial parts. APViT(Xue et al. 2022) proposes two Attentive Pooling modules with Vision Transformer to focus on the most discriminative features and discard the less relevant ones. POSTER V2(Mao et al. 2023), LA-Net(Wu and Cui 2023), and S2D(Chen et al. 2024) enable effective collaboration between facial landmark features and direct image features.
@@
Self-attention primarily captures dependencies within a sequence to focus on important information. SENet(Hu, Shen, and Sun 2018) introduced a plug-and-play channel attention module. BAM(Park et al. 2018) and CBAM(Woo et al. 2018) introduced spatial attention modules based on convolutions. However, these convolution-based spatial attention mechanisms can only perform local information interaction. The Non-Local Block(Wang et al. 2018) can capture long-range dependencies between any two positions in images. Subsequently, the Transformer(Vaswani et al. 2017) based on Scaled Dot-Product Attention became the state-of-the-art (SOTA) for natural language processing (NLP) tasks. ViT(Dosovitskiy et al. 2021) was the first to implement a pure transformer structure in the field of computer vision.
@@
Contrastive learning is a type of Self-Supervised Learning method(Chen et al. 2020; Chen, Xie, and He 2021). Supervised Contrastive Learning(Khosla et al. 2020) introduces labeled and different data of the same class for contrastive learning. Mutual Contrastive Learning(Chuanguang Yang 2022) collaboratively trains a cohort of models from the perspective of contrastive representation learning. Triplet loss(Schroff, Kalenichenko, and Philbin 2015) is a commonly used contrastive learning loss function in face recognition tasks. FEC(Vemulapalli and Agarwala 2019) learns a compact and continuous expression embedding space by comparing the similarity of three images using triplet loss.
@@
Intuitively, comparing the similarities and differences between images can make it easier to judge the category of an image. This similarity comes not only from high-level semantic features but also from relatively low-level fine-grained features. Based on the non-local block(Wang et al. 2018), we propose a spatial attention called Cross Similarity Attention (CSA) to focus on the similarity between different images. This is significantly different from the Dot-Product Attention in ViT(Dosovitskiy et al. 2021), and it does not require concatenation on spatial dimension of different images like the non-local block.
@@
Figure 3: Different types of attention on feature vectors h×w×c, where h=w=2, and c denotes channels. Blue and yellow represent features from different images, and each element corresponds to a 1×1×c vector. Double-arrowed lines denote the interaction values between corresponding two elements, and thicker lines indicate higher weights. The 4x4 green box in (d) represents the interaction matrix of (c).
Firstly, In standard ViT qkv self-attention (as shown in Fig. 3(a)), the query ${Q}$, key ${K}$, and value ${V}$ are all feature vectors derived from the same image. The output feature position is the same as the query feature position. The weight vectors can reflect the interactions between the input positional features and the output positional features.
However, if the query ${Q}$ comes from the features of one image p, while the key ${K}$ and values ${V}$ are derived from the features of a different image a, as shown in Fig. 3(b), some changes occur. The output positional features(a0, v0) in image ${a}$ are different from the query positional feature in image ${p}$(p0) (this is different from self-attention). The weight ${w}$ reflects the strength of the association with the values ${V}$, but it cannot reflect the strength of the association with the output vector out(t0) at the position in image ${a}$.
Therefore we consider only quantifying the importance of the positions in the queried image (key ${K}$), regardless of where the query ${Q}$ comes from, and using these weights to weight the features at the output positions (values ${V}$) of that image, as shown in Fig. 3(c) 3(d).
Figure 4: Interaction matrix of 7x7xC (H=7, W=7, channel=C) image features in the spatial dimension (HxW=49). (a) cosine similarity SA (b) dot product SA (c) euclidean distances SA (d) euclidean distances CA
@@
The original non-local block captures similarity among concatenated frames that is of practical significance, utilizing self-attention, but inflates the interaction matrix. In contrast, our CSA(Fig. 3(c)) efficiently computes similarity between different feature maps.
@@
We calculate the similarity interaction matrix using Euclidean distance as shown in Eq. 1 2, ${D_{cs}\left({Q,K}\right)_{ij}=\sqrt{\sum_{c=1}^{C}(Q_{ic}-K_{jc})^{2}}}$, ${Q,K,V\in\mathbb{R}^{l\times c}}$ where ${l=w\times h}$) and separately assess the importance of the queried positions in both images, thereby emphasizing the weights of common features with high similarity across different spatial locations within each image. Inspired by GCNet(Cao et al. 2019), which observed that attention maps calculated for different query positions are nearly identical, we aggregate the interaction matrix by rows or columns.
@@
Eq. 3 4 calculate the attention for image ${a}$. Here, ${\operatorname{norm}\left({S_{cs}\left({Q,K}\right)}_{i,:}\right)}$ normalizes the similarity matrix by row using the ${L2}$ norm to eliminate absolute differences in similarity calculated from query features at different positions, then sums the values for the same queried spatial position across columns. Finally, a mutually exclusive softmax function is employed to directly compute the weight distribution along the spatial dimension, where ${\sum_{j=1}^{l}\left(\operatorname{softmax}\left({S_{cs\_a}\left({Q,K}\right)}\right)^{T}\right)_{j}}=1$. Eq. 5 6 symmetrically calculate the attention for image ${p}$.
@@
By leveraging Cross Similarity Attention (Section 3), fine-grained features with high similarity between two images belonging to the same category can be easily extracted, and their weights enhanced. The resulting classifier, ${{Cls}_{cross}}$ , outperforms the backbone-based ${{Cls}_{base}}$ in discrimination. This approach is termed the Dual Cross Similarity (DCS) Network.
@@
To enhance similar features within the same category while emphasizing features with large distances between different categories. This is similar to the triplet loss(Schroff, Kalenichenko, and Philbin 2015), which reduces the feature differences within the same category and simultaneously increases the feature differences between different categories. However, if only the anchor-positive-negative triplet is introduced, the negative cannot construct a symmetric cross network with the anchor or positive due to the absence of a paired image from the same category. We have attempted to freeze the negative branch while only utilizing its redundant interfering features separated from the anchor branch, but there is a significant performance gap between this network and the joint training network, possibly due to the issue of gradient stability.
Therefore, we added another negative image, ingeniously forming the four branches into a centrally symmetric closed-loop network, as shown in Fig. 2. This approach allows us to calculate only two S matrices for the same category and two D matrices for different categories, with each interaction matrix being utilized by two adjacent images simultaneously.The features obtained from the backbone without the class token go through an FC layer to produce vectors ${QK}$ and ${V}$. When calculating the interaction matrices with the features of two adjacent images, the weights of ${QK}$ are shared and act as either ${Q}$ query or ${K}$ key. Consequently, the interaction matrices exhibit symmetry, where ${(Q2K1)^{T}=Q1K2}$ and ${(Q3K4)^{T}=Q4K3}$. When ${SD}$ fuses the matrices ${S}$ and ${D}$ by rows or columns, it reduces the need for two interactive calculations. The Quadruplet Cross Similarity(QCS) Network offers computational efficiency while achieving stability in multi-branch joint training.
@@
A straightforward approach is to introduce residual connections between ${{Cls}_{base}}$ and ${{Cls}_{cross}}$ for direct gradient feedback. We explored diverse connection methods, such as element-wise addition followed by Global Average Pooling (GAP) and Bilinear pooling (BP) to fuse and pool the feature vectors. Additionally, we considered treating ${{Cls}_{cross}}$ as the teacher model and ${{Cls}_{base}}$ as the student model, introducing a Kullback-Leibler (KL) divergence loss or L2 loss to construct a self-distillation learning(Zhang et al. 2019; Romero et al. 2015; Liu, Tang, and Wu 2020). We further integrated a ViT module into ${{Cls}_{cross}}$ to align with ${{Cls}_{base}}$, minimizing gradient conflicts on feature maps. We term the transfer of information from interactive contrastive learning in the cross module to the base network as Contrastive Residual Distillation.
@@
The joint training loss, utilizing cross entropy for all classifiers, is defined in Eq. 10. C set to 2 for DCS and 4 for QCS, The [] indicates an optional element.
@@
We adopted POSTER++ removing the landmark branch as the baseline and constructed our model upon it. All face images are resized to 224 × 224 pixels for training and testing. We use IR50(Deng et al. 2019), which is pretrained on the Ms-Celeb-1M dataset(Guo et al. 2016), as the backbone. We employ an Adam optimizer with Sharpness Awareness Minimization (SAM)(Foret et al. 2020) to train the model(200 epochs for RAF-DB, 100 epochs for FERPlus and 40 epochs for AffectNet). For the DCS, the batch size is set to 48, the learning rate is initialized as 9e-6 and an exponential decay learning rate scheduling with a gamma of 0.98 is employed. For the QCS, the batch size and the initial learning rate are reduced to around half. It is trained end-to-end on a single Nvidia RTX3090 GPU via Pytorch.
(a) row 4
(b) row 5
(c) row 6
(d) row 7
(e) D
(f) MSE
Figure 6: The training loss curves of ${Cls_{base}}$ and ${Cls_{cross}}$. (a)(b)(c)(d) correspond to the models in rows 4-7 of Table 1, respectively. (e)(f) correspond to the QCS, displaying only the anchor and pos branches.
@@
Initially, we trained a classification network utilizing the IR50+ViT backbone as a baseline. Then, a DCS Network with shared weight branches was constructed on the backbone to evaluate the performance impact of each component in our method through ablation studies on RAF-DB and FERPlus. Each experiment is repeated five times, then the maximum accuracy is reported. Unless otherwise stated, the experimental settings adopt the default configurations outlined above.
@@
Row 1 of Table 1 presents the accuracy of the baseline, while rows 3 and 4 show that our CSA outperforms SDPA in the comparison. For a fair comparison, all structures remain consistent except for varying core interaction approaches, with Multi-Head Attention and class token removed for maximum structural simplicity. Features (7×7×c) extracted from 3 abstraction levels of the backbone interact individually and are concatenated on the spatial dimension. CSA avoids computing similarities between features of different levels, and SDPA adopts the same approach. The FC and MLP (Multi-Layer Perceptron) of the interaction module share weights across the interaction branches.
@@
In Eq. 10, one of $\lambda_{1}$ and $\lambda_{2}$ is set to 1, while the other is set to 0. As shown in Table 1(row 7) and Fig. 6(d), adding the KL divergence loss narrows the losses gap between ${Cls_{base}}$ and ${Cls_{cross}}$, but the accuracy of the model decreases. This phenomenon becomes even more evident in the more complex QCS network using MSE loss, as shown in Fig. 6(f). It is possible that the cross module exhibits overfitting, influencing the base model’s generalization capacity. We set the dropout ratio ranging from 0.2 to 0.5 in the ViT modules to mitigate overfitting during training on AffectNet.
@@
Table 2: Accuracy (%) comparison of different cross matrices.
Table 2 shows S contributes more to improving the model performance compared to D. As can be seen from Fig. 6(e), ${Cls_{cross}}$, which is enhanced by D, does not show loss reduction compared to ${Cls_{base}}$. Fig. 5 shows that the inter-class similarity matrix has more dark blue areas (i.e., more bright yellow highlights in the inter-class distance matrix), indicating more dispersed attention. We infer that although weakening the features with small distances (high similarity) across categories helps separate redundant features, the simultaneously enhanced features with large distances do not necessarily contribute to classification. The SD matrix, resulting from the adaptive fusion of S and D, harnesses the strengths of both S and D to extract more clearly features.

### v2

@@ Abstract
Facial expression recognition faces challenges where labeled significant features in datasets are mixed with unlabeled redundant ones. In this paper, we introduce Cross Similarity Attention (CSA) to mine richer intrinsic information from image pairs, overcoming a limitation when the Scaled Dot-Product Attention of ViT is directly applied to calculate the similarity between two different images. Based on CSA, we simultaneously minimize intra-class differences and maximize inter-class differences at the fine-grained feature level through interactions among multiple branches. Contrastive residual distillation is utilized to transfer the information learned in the cross module back to the base network. We ingeniously design a four-branch centrally symmetric network, named Quadruplet Cross Similarity (QCS), which alleviates gradient conflicts arising from the cross module and achieves balanced and stable training. It can adaptively extract discriminative features while isolating redundant ones. The cross-attention modules exist during training, and only one base branch is retained during inference, resulting in no increase in inference time. Our proposed method achieves state-of-the-art performance on several FER datasets.
@@ Introduction
However, FER is still a complex task facing many difficulties such as Inter-class similarity and intra-class variances. Faces with varying expressions retain the common features that constitute a human face, yet same expressions can vary widely in pose, identity, illumination, etc. Traditional object classification methods struggle with complex face datasets due to assumptions of labeled feature dominance. Dividing datasets by labels for training can result in the learning of unlabeled redundant features if they’re non-randomly distributed, impacting classification accuracy. Fig. 1 illustrates one such impact. In reality, interference features are diverse, numerous, and unpredictable, exerting varying degrees of influence.
Figure 1: We often assume labeled features exhibit significant and dominant distributions within the dataset; however, unlabeled redundant features may impact this. Feature similarity can be leveraged to distinguish target features from redundant ones, based on whether they originate from the same label group or different label groups.
@@
How to extract discriminative features and how to eliminate the interference of redundant features are the core issues of FER. Using self-attention, some deep learning methods (Wang et al. 2020; Xue, Wang, and Guo 2021; Xue et al. 2022; Ma, Sun, and Li 2021) focus on key features, enhancing model discrimination. However, model performance hinges on the high-quality of training data. Some methods (Xie, Hu, and Wu 2019; Zhang et al. 2021; Cai et al. 2021; Ruan et al. 2021, 2022) study separating redundant from expression features using GANs or transformers. They extract expression information via decomposition and reconstruction, but often need redundant feature labels, such as age, gender, race, or rely on identity recognition modules. Additionally, interference decoupling FER (Li, Deng, and Du 2017) relies on one or several interference feature labels and can only address limited and specific interference factors. However, interference features are often unclear in type and quantity, and these methods that rely on labeled data often fail to achieve good performance when the types, quantities, and quality of labeled data are limited. Furthermore, when the expression feature labels are ambiguous (She et al. 2021; Zhang, Wang, and Deng 2021), it is unrealistic to clearly label all interference features.
Inspired by self-supervised contrastive learning methods(Yang et al. 2022; Gao and Patras 2024) that extract feature representations from unlabeled data, we attempt to conduct similarity comparisons between image pairs of different classes in order to isolate these redundant and interfering features, even though they are unlabeled. Additionally, we harness the similarity of same-class image pairs to extract more discriminative common features. Our method differs from those (Gao et al. 2020; Zhuang, Wang, and Qiao 2020) that utilize channel similarity and only perform comparisons between two images. It is also distinct from approaches(Yang et al. 2022; Vemulapalli and Agarwala 2019) that solely contrast high-level features among triplets.
Firstly, we propose Cross Similarity Attention (CSA), specifically tailored for mining fine-grained feature similarity between different images. We have identified a limitation when the Scaled Dot-Product Attention (SDPA) from ViT (Dosovitskiy et al. 2021) is directly applied to calculate image similarity: the similarity weights derived from the correlation between Q (query) and K (key) directly act on the feature vectors of the paired image, rather than the output feature vectors of the branch, resulting in attention that cannot obtain direct feedback (See Fig. 3(b) for details). Our CSA focuses solely on the importance of fine-grained similarity for the K (key), and directs attention directly between the input and output feature vectors of one branch. To our knowledge, we are the first to implement global spatial attention in the spatial dimension, analogous to the method of SENet(Hu, Shen, and Sun 2018) in the channel dimension. Previous works, such as BAM(Park et al. 2018) and CBAM(Woo et al. 2018), were limited to capturing only local spatial attention. The self-attention in ViT does not output a complete attention map after applying the softmax function; instead, it outputs results for each individual feature point.
@@
By leveraging CSA for the interaction module, we simultaneously minimize intra-class feature differences and maximize inter-class differences, mimicking the effect of Triplet loss (Schroff, Kalenichenko, and Philbin 2015) at the fine-grained feature level. In this way, our method enhances discriminative features and separates redundant ones. To effectively integrate the refined information from cross modules, we introduce residual connections between the cross classifier and the base classifier. To avoid gradient conflicts arising from multiple interaction modules across the anchor, positive, and negative branches, we introduce an additional image of the same category as the negative to construct a four-branch centrally symmetric closed-loop network. This framework reduces the computation of two interaction modules and enables a stable joint training. During inference, the multi-branch interaction modules are removed, leaving only a single base classifier, without increasing the inference time.
@@
  • We introduce Cross Similarity Attention (CSA), a global spatial attention mechanism specifically designed to mine fine-grained feature similarities across different images. It addresses the issue of indirect attention between input and output vector of one branch when computing similarity within the cross attention of ViT.
  • We propose Quadruplet Cross Similarity (QCS), a novelty four-branch, centrally symmetric closed-loop joint training framework based on CSA. It refines features by adaptively mining discriminative features within the same class while simultaneously separating interfering features across different classes.
  • Compared to methods that introduce additional landmark information, our method achieves state-of-the-art performance on several FER datasets by mining richer intrinsic information.
Figure 2: The framework of Quadruplet Cross Similarity Network. Joint training is performed on 4 classifiers ${{Cls}_{base}}$ based on a weight-shared backbone and 4 classifiers ${{Cls}_{cross}}$ enhanced by cross module, with only the red branch retained during inference. Anchor and pos are in the same category, so are neg and neg2. The matrix ${SD}$ performs attention refinement on matrix ${S}$ and matrix ${D}$ by rows or columns.
@@
Transfer(Xue, Wang, and Guo 2021) is the first attempt to apply Vision Transformer (ViT) to adaptively characterize the relations between different facial parts. APViT(Xue et al. 2022) proposes two Attentive Pooling modules with Vision Transformer to focus on the most discriminative features and discard the less relevant ones. POSTER(Zheng, Mendieta, and Chen 2023), POSTER V2(Mao et al. 2023), LA-Net(Wu and Cui 2023), and S2D(Chen et al. 2024) utilize ViT cross-fusion to integrate information from a facial landmark detector MobileFaceNet(Chen 2021), while FMAE(Ning, Salah, and Ertugrul 2024) is pretrained on a dataset of 9 million facial images (termed Face9M) for tasks related to AU detection and FER.
@@
Self-attention primarily captures dependencies within a sequence to focus on important information. SENet(Hu, Shen, and Sun 2018) introduced a plug-and-play channel attention module. BAM(Park et al. 2018) and CBAM(Woo et al. 2018) introduced spatial attention modules based on convolutions. However, these convolution-based spatial attention mechanisms can only perform local information interaction. The Non-Local Block(Wang et al. 2018) can capture long-range dependencies between any two positions in images. Subsequently, the Transformer(Vaswani et al. 2017) based on Scaled Dot-Product Attention became the state-of-the-art for natural language processing (NLP) tasks. ViT(Dosovitskiy et al. 2021) was the first to implement a pure transformer structure in the field of computer vision.
@@
Contrastive learning is a type of Self-Supervised Learning method(Chen et al. 2020; Chen, Xie, and He 2021). Supervised Contrastive Learning(Khosla et al. 2020) introduces labeled and different data of the same class for contrastive learning, improving the discriminative ability of the model. Triplet loss(Schroff, Kalenichenko, and Philbin 2015) is a commonly used contrastive learning loss function in face recognition tasks. FEC(Vemulapalli and Agarwala 2019) learns a compact and continuous expression embedding space by comparing the similarity of three images using Triplet loss.
@@
Intuitively, comparing the similarities and differences between images can make it easier to judge the category of an image. This similarity comes not only from high-level semantic features but also from relatively low-level fine-grained features. Based on the non-local block(Wang et al. 2018), we propose a spatial attention called Cross Similarity Attention (CSA) to focus on the similarity between different images. This is significantly different from the Scaled Dot-Product Attention in ViT(Dosovitskiy et al. 2021), and it does not require concat on spatial dimension of different images like the non-local block.
@@
[…]
Firstly, In standard ViT qkv self-attention (as shown in Fig. 3(a)), the query ${Q}$, key ${K}$, and value ${V}$ are all feature vectors derived from the same image. The attention weights acting on the queried vectors are equivalent to acting on the output vectors.
However, if the query ${Q}$ comes from the features of one image p, while the key ${K}$ and values ${V}$ are derived from the features of a different image a, as shown in Fig. 3(b), some changes occur. The output feature points in image ${a}$ are different from the query feature points in image ${p}$ which makes the output feature points unrelated to the calculation of the weights. The weight ${w}$ applies attention to the values ${V}$ and outputs to a feature point in the vector ${Out}$. The attention distribution within the vector ${Out}$ is difficult to assess through the weights.
Therefore we consider only quantifying the importance of the positions in the queried image (key ${K}$), regardless of where the query ${Q}$ comes from. Using these weights, we directly assign weight to all the features of values ${V}$ and output them to the vector ${Out}$ of the same branch, as shown in Fig. 3(c) 3(d).
(a) cosine similarity SA
(b) dot product SA
(c) euclidean distances SA
(d) euclidean distances CA
[…]
@@
Meanwhile, although the self-attention mechanism of the original ViT module can capture similarities among concatenated feature vectors, it also introduces interference from the similarities among the elements within the feature representation of a single image. The concat operation enlarges the interaction matrix. In contrast, our CSA (Figure 3(c)) computes similarities between different feature maps more efficiently.
@@
We calculate the similarity interaction matrix using Euclidean distance as shown in Eq. 1 2, ${D_{cs}\left({Q,K}\right)_{ij}=\sqrt{\sum_{c=1}^{C}(Q_{ic}-K_{jc})^{2}}}$, ${Q,K,V\in\mathbb{R}^{l\times c}}$ where ${l=h\times w}$) and separately assess the importance of the queried positions in both images, thereby emphasizing the weights of common features with high similarity across different spatial locations within each image. Inspired by GCNet(Cao et al. 2019), which observed that attention maps calculated for different query positions are nearly identical, we aggregate the interaction matrix by rows or columns.
@@
Eq. 3 4 calculate the attention for image ${a}$. Here, ${\operatorname{norm}\left({S_{cs}\left({Q,K}\right)}_{i,:}\right)}$ normalizes the similarity matrix by row using the ${L2}$ norm to eliminate absolute differences in similarity calculated from query features at different positions, then sums the values for the same queried spatial position across columns. Finally, a mutually exclusive softmax function is employed to directly compute the weight distribution along the spatial dimension, where ${\sum_{j=1}^{l}\left(\operatorname{softmax}\left({S_{cs\_a}\left({Q,K}\right)}\right)^{T}\right)_{j}}=1$. We use softmax in the global spatial dimension because it operates on the similarity map of two different images, rather than on the feature map of a single image. Thus, there is no concern that applying Eq. 4 will disrupt the semantic structure of the image. Eq. 5 6 symmetrically calculate the attention for image ${p}$.
@@
By referencing the method of interactive comparison between two images (Gao et al. 2020; Zhuang, Wang, and Qiao 2020), we utilize Cross Similarity Attention to enhance fine-grained features that exhibit high similarity between two images belonging to the same category. The resulting classifier, ${{Cls}_{cross}}$ , outperforms the backbone-based ${{Cls}_{base}}$ in discrimination. This approach is termed the Dual Cross Similarity (DCS) Network (Fig. 5).
Figure 5: The framework of Dual Cross Similarity Network.
@@
The Triplet loss(Schroff, Kalenichenko, and Philbin 2015) relies on comparing distances between embedding representations. Our method utilizes CSA in conjunction with the supervised classifier ${Cls_{cross}}$ to achieve two objectives at the fine-grained feature level: it brings the distances between images of the same class closer (enhancing fine-grained features with high similarity) and pushes the distances between images of different classes farther away (weakening fine-grained features with high similarity, i.e., small distances).
The backbone of the anchor and pos branches are directly connected to the ${Cls_{base}}$ for supervised learning. If the classifier is not added to the neg branch, gradient conflicts will arise among the anchor, pos, and neg branches. Considering that the cross similarity modules S and D learn different content, the neg branch cannot be frozen during training (See the Appendix). Through various experimental attempts, it has been difficult to train effective models using networks constructed with triplet.
To avoid gradient conflicts, the pos, anchor, and neg branches need to have equal status throughout the network, with each branch connected to the supervised classifier ${Cls_{base}}$. Ultimately, we ingeniously added another negative image, forming the four branches into a centrally symmetric closed-loop network called Quadruplet Cross Similarity(QCS) Network, thereby avoiding gradient conflicts and enhancing training stability, as shown in Fig. 2.
Meanwhile, the QCS offers computational efficiency. It allows us to calculate only two S matrices for the same category and two D matrices for different categories, with each interaction matrix utilized by two adjacent images simultaneously. The features obtained from the backbone without the class token go through an FC layer to produce vectors ${QK}$ and ${V}$. When calculating the interaction matrices with the features of two adjacent images, the weights of ${QK}$ are shared and act as either query ${Q}$ or key ${K}$. Consequently, the interaction matrices exhibit symmetry, where ${(Q2K1)^{T}=Q1K2}$ and ${(Q3K4)^{T}=Q4K3}$. When ${SD}$ fuses the matrices ${S}$ and ${D}$ by rows or columns, it reduces the need for two interactive calculations.
@@
A straightforward approach is to introduce residual connections between ${{Cls}_{base}}$ and ${{Cls}_{cross}}$ for direct gradient feedback. We explored diverse connection methods, such as element-wise addition followed by Global Average Pooling (GAP) and Bilinear pooling (BP) to fuse and pool the feature vectors. Additionally, we considered treating ${{Cls}_{cross}}$ as the teacher model and ${{Cls}_{base}}$ as the student model, introducing a Kullback-Leibler (KL) divergence loss or L2 loss to construct a self-distillation learning(Zhang et al. 2019; Liu, Tang, and Wu 2020). We further integrated a ViT module into ${{Cls}_{cross}}$ to align with ${{Cls}_{base}}$, minimizing gradient conflicts on feature maps. We term the transfer of information from interactive contrastive learning in the cross module to the base network as Contrastive Residual Distillation.
@@
The joint training loss, utilizing cross entropy for all classifiers, is defined in Eq. 10. C is set to 2 for DCS and 4 for QCS, $\lambda_{1}$ is set to 1, The [] indicates an optional element.
@@
[…]
(a) row 6
(b) row 7
(c) MSE
(d) D matrix
Figure 7: The training loss curves of ${Cls_{base}}$ and ${Cls_{cross}}$. (a)(b) correspond to the models in rows 6-7 of Table 1, respectively. (c)(d) correspond to the QCS, displaying only the anchor and pos branches.
@@
Initially, we trained a classification network utilizing the IR50+ViT backbone as a baseline. Row 1 of Table 1 presents the accuracy of the baseline. Then, the DCS Network with shared weight branches was constructed on the backbone to evaluate the performance impact of each component in our method through ablation studies on RAF-DB and FERPlus.
@@
[…]
@@
In Eq. 10, one of $\lambda_{2}$ and $\lambda_{3}$ is set to 1, while the other is set to 0. As shown in Table 1(row 7) and Fig. 7(b), adding the KL divergence loss narrows the losses gap between ${Cls_{base}}$ and ${Cls_{cross}}$, but the accuracy of the model decreases. This phenomenon becomes even more evident in the more complex QCS network using MSE loss, as shown in Fig. 7(c). It is possible that the cross module exhibits overfitting, influencing the base model’s generalization capacity. We set the dropout ratio ranging from 0.1 to 0.4 in the ViT modules to mitigate overfitting during training on AffectNet.
@@
Table 2: Comparison(%) of different cross matrices.
Table 2 shows S contributes more to improving the model performance compared to D. As can be seen from Fig. 7(d), ${Cls_{cross}}$, which is enhanced by D, does not show loss reduction compared to ${Cls_{base}}$. Fig. 6 shows that the inter-class similarity matrix has more dark blue areas (i.e., more bright yellow highlights in the inter-class distance matrix), indicating more dispersed attention. We infer that although weakening the features with small distances (high similarity) across categories helps separate redundant features, the simultaneously enhanced features with large distances do not necessarily contribute to classification. The SD matrix, resulting from the adaptive fusion of S and D, harnesses the strengths of both S and D to extract more clearly features.